Can you help me-- Solve the system by the addition method. Determine whether the equations are independent, dependent or inconsistent.
4x+3y = 2
-12x – 9 y = - 6
Thank you all

Answer by stanbon(75887)

Solve the system by the addition method. Determine whether the equations are independent, dependent or inconsistent.
4x+3y = 2
-12x – 9 y = - 6
------------------------
Multiply thru the 1st equation by 3 to get:
12x + 9y = 6
-12x- 9y = -6
---------
Addiing you get:
0 = 0
=================
That means the two equations have the same graph.
The system is said to be "dependent".
The solution set is the set of points of the line.
==========================================
Cheers,
Stan H.